OpenStudy (anonymous):

need it urgently.....plse help me to understand the question and plsee solve it

OpenStudy (amistre64):

it would help to know the slope (tangent) of the given degrees

OpenStudy (amistre64):

given a slope (m) and a point (a,b); a line can be constructed as: y = m(x-a) + b

OpenStudy (amistre64):

assuming we can find m, how would you set up the 4 lines

OpenStudy (anonymous):

srry 'i don't know

OpenStudy (amistre64):

i posted the instructions, the format, the setup .... you will have to tell me what you do not understand about it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

u wrote set up 4 lines..what does that mean

OpenStudy (amistre64):

the question is asking you for 4 lines ....given a slope (m) and a point (a,b); a line can be constructed as: y = m(x-a) + b assuming we can determine the slope (m) later, can you set up the 4 lines that the question is asking for?

OpenStudy (anonymous):

but if we want to set up 4 line, we need to put a and b value in y = m(x-a)+b

OpenStudy (amistre64):

correct, can you give me the points that they want? one is directly stated, and the other is infered.

OpenStudy (anonymous):

a=0 and b= 2

OpenStudy (anonymous):

second one is a=0, b=0 third one is a=0, b=2

OpenStudy (amistre64):

(0,2) is one of them y = mx + 2 will take care of 2 lines thru (0,2) ----------------------------------------- the point that is 2 below the origin is: (0,-2) y = mx - 2 will take care of 2 lines thru (0,-2)

OpenStudy (amistre64):

itll look like this 

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

so, we need to determine the slope that will work, do you have an idea on how we could determine those slopes?

OpenStudy (anonymous):

hmm.no

OpenStudy (amistre64):

the give us a measures of angles in degrees. to me this means that we can use trigonometry to determine the slope of the lines: m = y/x tan(angle) = y/x

OpenStudy (amistre64):

they give us 60 and 120 degrees; m1=tan(60) and m2 = tan(120) y1 = tan(60)x + 2 y2 = tan(120)x + 2 y3 = tan(60)x - 2 y4 = tan(120)x - 2

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

just as an aside: tan(60) = sqrt(3) tan(120) = -sqrt(3)

OpenStudy (anonymous):

so what after this

OpenStudy (anonymous):

@amistre64 what to do next

OpenStudy (amistre64):

well, seeing how we have "found" the 4 lines .... or at least equations for them; unless there is a picture to choose from id say its done

OpenStudy (anonymous):

it means only in this question we have to find y1, y2, y3 and y4 that's it

OpenStudy (amistre64):

ok, since we found them ... there is nothing left to do with the question :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@amistre64 thank u

OpenStudy (amistre64):

good luck